OpenStudy (thedj4jc):

Simplify 18 - 2[x + (x - 5)]. 8 - 4x 28 - 4x 28 - 2x

OpenStudy (anonymous):

4 (7-x)?

OpenStudy (anonymous):

18-2[2x-5] =18-4x+10 =-4x+28

OpenStudy (thedj4jc):

Thank you!!!